Question 2 Two parallel plates, each charged equally and oppositely to the other, are separated by 1.5500 cm. A proton is let go from rest at the positive plates surface and, at the same time, an electron is let go from rest at the negative plates surface. What is the distance between the negative plate and the point where the proton and the electron go by each other? Note: unlike most questions, this one will need your answer correct to 5 significant digits. Make sure you enter that many into the box below.
Answer #1

From newtons 2nd law

F =m*a

a =F /m

From kinematic equation

s=u*t+1/2*a*t^2

Let d be the distance from the negative plate.

de me rmp -2구 い550CM-d 9-11X10